# Proprioceptive State Estimation for Amphibious Tactile Sensing

### v1

@@ 3.2 Geometry-Related Deformation Energy Function
where $\mathbf{x}\in{\mathbb{R}^{3\times{n}}}$ contains all the discretized vertices location of the soft body $\mathcal{M}$.
3.3 Rigidity-Aware Aggregated Multi-handles Constraints from Visual Observation
Compared to multi-view systems, the wide availability and ease of use make monocular cameras one of the best choices to achieve environmental perception. However, deformable shape reconstruction from 2D image observations is well-known as an ill-posed inverse problem and has been actively researched [54]. To make this problem trackable, we leverage the volumetric discretized model described above and introduce rigidity-aware aggregated multi-handle constraints to reliably reconstruct the soft finger’s deformed shape.
As shown in Figs. 2A(iii)&(iv), a rigid plate is mechanically coupled with part of soft material, resulting in a uniform rigid transformation $g$ for each attached node in the discrete model $M$.
@@ 3.3 Rigidity-Aware Aggregated Multi-handles Constraints from Visual Observation
where $\mathbf{x}_{h}\in{\mathbb{R}^{3\times{p}}}$ contains deformed location of $p$ vertices related to the rigidity-aware aggregated multi-handles constraints, while $\mathbf{X}_{h}\in{\mathbb{R}^{3\times{p}}}$ contains the corresponded undeformed vertices location. The rigid transformation $g$ is estimated using fiducial markers, widely used in robot vision tasks.
3.4 Shape Estimation by Geometric Optimization
With the discrete energy functional Eq. 7 of the given soft body $\mathcal{M}$ and observed kinematics constraints Eq. 8, soft body shape estimation can be directly translated into a constrained geometry optimization problem as:
$\min_{\mathbf{x}}\quad$ $\sum_{t_{j}\in{\mathcal{T}}}{\Psi{(\mathbf{A}_{t_{j}}(\mathbf{x}))}},$ (9)
@@ 3.4 Shape Estimation by Geometric Optimization
Instead of considering kinematics constraints as hard boundary conditions, we enforce them by appending quadratic penalty functions to original functional $E(\mathbf{x})$ in Eq. 7 for easier handling, which results in
$\tilde{E}(\mathbf{x})=\sum_{t_{j}\in{\mathcal{T}}}{\Psi{(\mathbf{A}_{t_{j}}(\mathbf{x}))}+\omega||\mathbf{x}_{h}-g(\mathbf{X}_{h})||^{2}},$ (10)
and the deformed shape can be estimated by minimizing the augmented energy functional in Eq. 10 as:
@@
where $\omega$ is the penalty parameter for the corresponding unconstrained minimization problem. Note that a more considerable penalty weight will lead to better constraint satisfaction but poor numerical conditions.
In practice, we set $\omega=10^{5}$ and compute the deformed vertices positions $\mathcal{V}$ by iteratively minimizing Eq. 11 using a Newton-type solver as shown in Alg. 1. As shown in Fig. 2B, a series of physically plausible deformations of the soft finger under observed constraints are reconstructed in real-time using the proposed optimization approach.
@@
4 Object Shape Estimation using Tactile Sensing
@@
While proprioception refers to being aware of one’s movement, tactile sensing involves gathering information about the external environment through the sense of touch. This section presents an object shape estimation approach by extending the proprioceptive state estimation method proposed in Section 3 to tactile sensing.
@@
Since passive adaptive structures can quickly adapt their shape to conform to contact objects, we could infer shape-related contact information from the estimated shape of the soft finger during contact. By assuming that the contact patch on the soft finger coincides with that of the object during grasping, we can predict object surface topography using spatially distributed contact points on the contact interface.
4.1 Contact Interface Points Extraction
Based on the spatial discretization model in Section 3.1, an indexed set =$\mathcal{I}=\{c_{1},c_{2},...,c_{k}\}$ of nodes located at upper area of the soft finger mesh $\mathcal{M}$ are extracted as contact interface points, as shown in Fig. 3A.
@@ 4.1 Contact Interface Points Extraction
Figure 3: Pipeline for contact interface geometry sensing using deformed positions of soft finger mesh nodes. (A) Owing to the ability of the soft finger to deform and adapt its shape to fit the contours of the object being grasped, we take the deformed soft finger mesh nodes as approximate multi-contact points on the contact interface. (B) In addition to the mesh nodes $\mathbf{x}_{c}$ on the contact interface, auxiliary training points $\mathbf{x}^{-}_{c}$ and $\mathbf{x}^{+}_{c}$ are generated in this step to increase the accuracy of the implicit surface reconstruction. (C) Gaussian process implicit surface model is adopted for contact object surface patch estimation.
@@
Each time with the observed AMH constraints input, we could determine the positions of these contact interface points by first solving Eq. 11, then extracting corresponding nodes using indexed set $\mathcal{I}$, from solved deformed positions of vertices $\mathcal{V}$: $\mathbf{x}_{c}=\{\mathbf{x}_{i}|\mathbf{x}_{i}\in{\mathcal{V}},i\in{\mathcal{I}}\}$.
4.2 Implict Surface Representation for Object Shape
@@
Considering the grasping action using a soft finger as a multi-point tactile probing, the object surface patches could be progressively reconstructed by these gripping actions with collected positions of contact interface points $\mathbf{x}_{c}$ extracted from the soft finger.
@@
An implicit surface representation is defined by a function that can be evaluated at any point in space, which yields a value indicating whether the point is inside the object, outside the object, or on the object’s surface. For the 3-D space considered in our problem, this function $f$ is defined as:
@@
$f(\mathbf{x}):\mathbb{R}^{3}\rightarrow{\mathbb{R}}=\begin{cases}<0,&\mbox{ if$\mathbf{x}$ inside the object};\\ =0,&\mbox{ if $\mathbf{x}$ on the surface};\\ >0,&\mbox{ if $\mathbf{x}$ outside the object}.\end{cases}$ (12)
@@
As is shown in Fig. 3B, we only collected positions of partial contact interface points $\mathbf{x}_{c}$, which are assumed to coincide with the object surface for each gripping action. While surface points are observed, we do not explicitly observe off-surface or internal points exemplars. For those unobserved cases in Eq. 12, we generate control points of the corresponding two types to express the directional information of the surface using the method described in [55].
4.3 GPIS for Surface Estimation
@@
An object’s shape is estimated by finding the points with zero value of implicit surface function Eq. 12 (i.e., the isosurface) in a 3D region of interest. Gaussian process implicit surface (GPIS), as a non-parametric probabilistic method [56], is a powerful tool for object surface reconstruction from partial or noisy 3D data and often used in the context of tactile and haptic exploration [57, 58].
@@
A Gaussian Process (GP) is a collection of $N$ random variables with a joint Gaussian distribution and, therefore, can be completely specified by its mean and covariance functions. Given the collected contact interface points togeter with generated control points positions $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}\}$ for each gripping action and corresponding observed values $\mathcal{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$, such that $\mathbf{y}_{i}=f(\mathbf{x}_{i})+\epsilon$, where $\epsilon\sim{\mathcal{N}(0,\sigma^{2}_{\epsilon})}$ denotes Gaussian noise with zero mean and $\sigma^{2}_{\epsilon}$ variance, the GP can be written as $f(\mathbf{x})\sim{\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))}$, where, $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function [59]. In our implementation, we use the radial basis function kernel, which is characterized by the two hyper-parameters the variance $\sigma^{2}_{f}$ and the length scale $l$:
@@
5 Experimental Results
We implemented our geometric optimization-based shape estimation algorithm (Alg. 1) for the soft finger in C++ and tested it on a standard PC with an Intel Core i7 3.8GHz CPU and 16 GB RAM. With the help of algorithmic differentiation of the numerical solver, Eigen [60], our system can realize the computation of the soft finger shape deformation with 3k tetrahedra in real-time (up to 20 fps).
This section presents the results of proprioceptive shape estimation of the soft finger using two experiment setups. One experiment leverages motion capture markers as ground truth, which provides high-precision but sparse measurements. The other experiment uses a Touch haptic device for ground truth data collection, which is less accurate but contains larger measuring coverage on the soft finger.
Table 1: Run Time and Mean Error Comparisons of ABAQUS, ARAP, and Our Method
@@ 5 Experimental Results
  • $*{}^{*}$: The benchmark of mean error is ABAQUS.
5.1 Comparisons with the Conventional Method
This section presents the performance of our shape estimation method by comparing it with two conventional methods, ABAQUS and as-rigid-as-possible (ARAP) [61]. Table 1 compares run time and mean error of shape estimation. Five different meshes with different numbers of elements, including 1k, 1.5k, 3k, 6k, and 12k, were provided, and the three methods were tested on the meshes with six different kinds of motions applied on AMH shown in Fig. 2B. It can be observed that our method can be 40 times faster than ABAQUS and two times faster than ARAP. We also compared the mean errors of all nodes estimated by our method and ARAP when benchmarked against ABAQUS. The results show that our method’s mean error becomes more minor, from 0.346 mm to 0.086 mm, as the number of elements becomes more extensive, while ARAP’s error remains 0.7-1.0 mm for different meshes. We also observe that the error of our method significantly decreases from 1k to 1.5k mesh while the error reduction from 1.5k to 12k is minimal. In terms of both running time and error, our approach has a significant advantage over ABAQUS and ARAP. In addition, 1.5k mesh is the most appropriate for our method, with both faster run speed and smaller error, which was selected for real-time estimation in the following experiments.
Figure 4: Estimated marker deformation obtained by proposed proprioceptive shape estimation method. (A) Experimental setup, including the soft finger, embedded with an RGB camera, a manual three-axis motion test platform, and six motion capture markers $m_{1},m_{2},...,m_{6}$, rigidly attached to the soft finger. (B) The estimated position of the marker $x_{m_{k}}^{\prime}$ is calculated using the barycentric coordinate of the corresponding attached tetrahedron $t_{k}$, while the ground truth reading $x_{m_{k}}$ is obtained from the motion capture system. (C) The corresponding error for each marker’s three-dimensional deformation and total norm.
5.2 Deformation Estimation with Motion Capture Markers
@@
As is shown in Fig. 4A, the soft finger is mounted on a three-axis motion platform to perform a set of deformed configurations for shape estimation. The test platform is operated manually to generate a set of different contact configurations between the soft finger and the indenter. During the process, the embedded camera streams real-time image data at a resolution of 640 × 480 pixels. The detected AMH rigid motion, using an off-the-shelf ArUco detection library [62], is fed into our implemented program for shape estimation.
@@
Nokov vision system with client software tool tracks finger deformation and motion. A total of six motion capture markers with an 8 mm radius were used in the experiment. Markers are divided into three pairs and are rigidly attached to the fingertip ($m_{5},m_{6}$), the first layer ($m_{3},m_{4}$), and the second layer of the soft finger ($m_{1},m_{2}$), respectively. Besides the markers used for soft finger deformation tracking, three other markers are attached to the platform and used as the reference reading to align the motion capture system’s reference frame with the deformation computation coordinate frame.
@@
The markers are attached to the soft finger with rigid links. To avoid occlusion while tracking markers, we design the connecting links in three types of lengths, as shown in Fig. 4B. We assume that each marker is rigidly attached to the nearest tetrahedron on the parameterized mesh model $\mathcal{M}$ and represent the location of the estimated marker using barycentric coordinates of the corresponding tetrahedron element in the deformed configuration of the soft finger:
@@
Due to the rigid connection assumption, the barycentric coordinates $\bm{\lambda}_{t_{k}}$ are constant during the deformation. After solving barycentric coordinates in Eq. 16 using the initial undeformed vertex position of the tetrahedron and corresponding tracked marker position without contact, the marker position prediction model is a linear combination of deformed vertex position of the corresponding tetrahedron resulting from geometric optimization Alg. 1 using calibrated barycentric coefficients. See Movie S1 in the Supplementary Materials for a video demonstration.
@@
We visualize the error distribution with 3k data pairs of six markers’ estimated and ground truth positions as illustrated in Fig. 4C. The norm of the total error of all experiment records for six markers is around $(0,3)$ mm, while error distribution along each axis is centered around the $(-2,2)$ mm range. As the marker prediction model in Eq. 16 comprises calibration and geometric optimization, the error distribution of six sparse markers may not be enough to validate the proposed method, leading to another dense measurement experiment in the following section.
5.3 Deformation Estimation using Touch Haptic Device
To further evaluate the proprioceptive shape estimation method using dense measurements, we designed another validation experiment with the position of pen-nib of a touch haptic device as ground truth measurement.

### v2

@@ 3.2.2 Geometry-Related Deformation Energy Function
where $\mathbf{x}\in{\mathbb{R}^{3\times{n}}}$ contains all discretized vertices location of the soft body $\mathcal{M}$.
3.2.3 Rigidity-Aware Aggregated Multi-Handles Constraints
Monocular cameras are generally considered the primary sensory for environmental perception due to their ease of use and availability, compared to multi-view systems. However, deformable shape reconstruction from 2D image observations is well-known as an ill-posed inverse problem and has been actively researched [63]. We leverage the proposed volumetric discretized model and introduce rigidity-aware Aggregated Multi-Handle (AMH) constraints to make this problem trackable, aiming at reconstructing the soft finger’s deformed shape reliably.
We model the mechanical coupling of the rigid plate for the fiducial marker in Fig. 1A as a uniform rigid transformation $g$ for each attached node in the discrete model $M$, as shown in Figs. 2A(iii)&(iv):
@@ 3.2.3 Rigidity-Aware Aggregated Multi-Handles Constraints
where $\mathbf{x}_{h}\in{\mathbb{R}^{3\times{p}}}$ contains deformed location of $p$ vertices related to the rigidity-aware AMH constraints while $\mathbf{X}_{h}\in{\mathbb{R}^{3\times{p}}}$ contains the corresponding undeformed vertices location. The rigid transformation $g$ is estimated by fiducial markers widely used in robotic vision.
3.2.4 Geometric Optimization for Shape Estimation
With the discrete energy function Eq. (7) of the given soft body $\mathcal{M}$ and observed kinematics constraints Eq. (8), soft body shape estimation can be directly translated into a constrained geometry optimization problem as:
$\min_{\mathbf{x}}$ $\sum_{t_{j}\in{\mathcal{T}}}{\Psi{(\mathbf{A}_{t_{j}}(\mathbf{x}))}},$ (9)
@@ 3.2.4 Geometric Optimization for Shape Estimation
Instead of considering kinematics constraints as hard boundary conditions, we enforce them by appending quadratic penalty terms to $E(\mathbf{x})$ in Eq. (7) for easier handling, which results in
$\tilde{E}(\mathbf{x})=\sum_{t_{j}\in{\mathcal{T}}}{\Psi{(\mathbf{A}_{t_{j}}(\mathbf{x}))}+\omega||\mathbf{x}_{h}-g(\mathbf{X}_{h})||^{2}}.$ (10)
As illustrated in Fig. 2A(v), we can achieve deformed shape estimation by minimizing the augmented energy function in Eq. (10) as:
@@
where $\omega$ is the penalty parameter for the corresponding unconstrained minimization problem. A greater penalty weight will lead to better constraint satisfaction but poorer numerical conditions.
In practice, we set $\omega=10^{5}$ and compute the deformed vertices positions $\mathcal{V}$ by iteratively minimizing Eq. (11) using a Newton-type solver shown in Alg. 1. As shown in Fig. 2B, a series of physically plausible deformations of the soft finger under observed constraints are reconstructed in real time using our proposed optimization approach.
@@
3.3 Object Shape Estimation using Tactile Sensing
@@
While proprioception refers to being aware of one’s movement, tactile sensing involves gathering information about the external environment through the sense of touch. This section presents an object shape estimation approach by extending the PropSE method proposed in the previous section to tactile sensing.
@@
Since our soft finger can provide large-scale, adaptive deformation conforming to the object’s geometric features through contact, we could infer shape-related contact information from the finger’s estimated shape during the process. We assume the soft finger’s contact patch coincides with that of the object during grasping. As a result, we can predict object surface topography using spatially distributed contact points on the touching interface.
@@
Figure 3: Pipeline for contact interface geometry sensing using deformed positions of soft finger mesh nodes. (A) Because the soft finger can deform and adapt its shape to fit the contours of the object being grasped, we take the deformed soft finger mesh nodes as approximate multi-contact points on the contact interface. (B) In addition to the mesh nodes $\mathbf{x}_{c}$ on the contact interface, auxiliary training points $\mathbf{x}^{-}_{c}$ and $\mathbf{x}^{+}_{c}$ are generated in this step to increase the accuracy of the implicit surface reconstruction. (C) Gaussian process implicit surface model is adopted for contact object surface patch estimation.
3.3.1 Contact Interface Points Extraction
Based on the spatial discretization model in Section 3.2.1, an indexed set =$\mathcal{I}=\{c_{1},c_{2},...,c_{k}\}$ of nodes located at the upper area of the soft finger mesh $\mathcal{M}$ are extracted as contact interface points, as shown in Fig. 3A.
@@ 3.3.1 Contact Interface Points Extraction
With each of the observed AMH constraints input, we could determine the positions of these contact interface points by first solving Eq. (11), then extracting corresponding nodes using indexed set $\mathcal{I}$ by solving the deformed positions of vertices $\mathcal{V}$: $\mathbf{x}_{c}=\{\mathbf{x}_{i}|\mathbf{x}_{i}\in{\mathcal{V}},i\in{\mathcal{I}}\}$.
3.3.2 Implict Surface Representation for Object Shape
@@
Considering the grasping action using a soft finger as a multi-point tactile probe, the object surface patches could be progressively reconstructed by these gripping actions with collected positions of contact interface points $\mathbf{x}_{c}$ extracted from the soft finger.
@@
An implicit surface representation is defined by a function that can be evaluated at any point in space, yielding a value indicating whether the point is inside the object, outside the object, or on the object’s surface. For the 3-D space considered in our problem, this function $f:\mathbb{R}^{3}\rightarrow{\mathbb{R}}$ is defined as:
@@
$f(\mathbf{x})\begin{cases}<0,&\mbox{ if $\mathbf{x}$ inside the object};\\ =0,&\mbox{ if $\mathbf{x}$ on the surface};\\ >0,&\mbox{ if $\mathbf{x}$ outside the object}.\end{cases}$ (12)
@@
As is shown in Fig. 3B, we only collected positions of partial contact interface points $\mathbf{x}_{c}$, which are assumed to coincide with the object surface for each gripping action. While surface points are observed, we do not explicitly observe off-surface or internal points exemplars. For those unobserved cases in Eq. (12), we generate control points of the corresponding two types to express the directional information of the surface using the method described in [64].
3.3.3 GPIS for Surface Estimation
@@
An object’s shape is estimated by finding the points with zero value of implicit surface function Eq. (12) (i.e., the isosurface) in the 3D region of interest. The Gaussian Process Implicit Surface (GPIS) method can be used as a tool for object surface reconstruction from partial or noisy 3D data. It is a non-parametric probabilistic method often used for tactile and haptic exploration [65, 66].
@@
A Gaussian Process (GP) is a collection of $N$ random variables with a joint Gaussian distribution which can be specified using its mean and covariance functions. The collected contact interface point and the generated control point positions $\mathcal{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}\}$ for each grasping action and the corresponding observed values are denoted as $\mathcal{Y}=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$. Here, $\mathbf{y}_{i}=f(\mathbf{x}_{i})+\epsilon$, where $\epsilon\sim{\mathcal{N}(0,\sigma^{2}_{\epsilon})}$ denotes Gaussian noise with zero mean and $\sigma^{2}_{\epsilon}$ variance. As a result, the GP can be written as $f(\mathbf{x})\sim{\mathcal{GP}(m(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime}))}$, where $m(\mathbf{x})$ is the mean function and $k(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function [67].
In our implementation, we used the radial basis function kernel, which is characterized by the two hyper-parameters, the variance $\sigma^{2}_{f}$ and the length scale $l$, expressed as the following:
@@
4 Results
4.1 On Vision-based Proprioceptive State Estimation
Here, we first present the benchmarking results against two widely adopted methods to demonstrate the superior performance of our proposed vision-based PropSE method. Then, we present the results of our proposed vision-based PropSE method using two experiment setups. One leverages motion capture markers as ground truth, providing high-precision but sparse measurements. The other uses a touch-haptic device for ground truth data collection, which is less accurate but contains larger measuring coverage on the soft finger.
The implementation of the proposed geometric optimization-based algorithm (Alg. 1) was developed in C++ and evaluated on a computer with an Intel Core i7 3.8 GHz CPU and 16 GB of RAM. By leveraging the capabilities of algorithmic differentiation within the numerical solver, Eigen [68], this system demonstrated the capability to compute deformations involving 1,500 tetrahedra in real-time, achieving frame rates up to 20 fps.
4.1.1 Comparison with the Conventional Methods
We performed a comparative analysis with two widely adopted techniques to showcase the efficacy of our shape estimation method. One is Abaqus, a premier finite element analysis (FEA) software extensively applied in structural analysis and deformation modeling across various engineering disciplines. This comparison aims to highlight the versatility and precision of our approach within contexts requiring intricate modeling capabilities. (Please refer to Appendix A for further details concerning the Abaqus simulation.)
The other is the As-Rigid-As-Possible (ARAP) method [69], a widely adopted method in digital geometry processing for estimating object shapes through minimal rigid deformation. This comparison is particularly valuable, as ARAP’s principles of shape preservation align closely with the core objectives of our shape estimation task, providing valuable benchmarking. (Please refer to Appendix B for further details regarding our implementation.)
Table 1 compares our proposed method’s run time and mean error with those mentioned earlier. Each method is evaluated on five meshes with increasing resolutions, resulting in 1k, 1.5k, 3k, 6k, and 12k elements. The soft finger underwent six motions applied on the AMH shown in Fig. 2B with all the deformation data recorded. We treat the results from Abaqus as the ground truth. Results show that our method is 40 to 700 times faster than Abaqus and 1 to 2 times faster than ARAP at different resolutions. We also compared the mean errors of all nodes estimated by our method and ARAP when benchmarked against Abaqus. The results show that our method’s mean error decreases significantly, from 0.346 mm to 0.086 mm, as the number of elements increases. The ARAP’s error ranges from 0.7 mm to 1.0 mm for different meshes. Our approach shows significant advantages over Abaqus and ARAP regarding running time and accuracy.
Table 1: Run Time and Mean Error Comparisons of Abaqus, ARAP, and Our Method.
@@ 4.1.1 Comparison with the Conventional Methods
  • : The benchmark of mean error is Abaqus.
The optimization solver deployed to minimize the ARAP energy leverages the local/global method (as detailed in Appendix B). While this solver efficiently approximates the local minimum, its approach to convergence towards a numerical minimum necessitates a considerable number of iterations, a characteristic underscored during implementation [62]. We fixed the number of iterations at 10 for our benchmarking procedure to achieve convergence. This predefined iteration limit could account for the observed comparative slowness of the ARAP optimization solver relative to our proposed method. Regarding the evaluation of mean error, the suboptimal performance of ARAP, as compared to ours, might be attributed to the local/global optimization solver settings. Moreover, the deformation energy model used by ARAP might not fully encompass the non-linear deformation behaviors of our soft robotic fingers.
We also observe that the error of our method decreases most dramatically when the number of elements increases from 1k to 1.5k, and the error reduction from 1.5k to 6k is marginal. Hence, the mesh with 1.5k elements is the most appropriate for our method, achieving both faster run speed and minor error, which was selected for real-time estimation in the following experiments. (Please refer to Appendix C for additional results on Algorithm 1 parameter.)
4.1.2 Deformation Estimation with Motion Capture Markers
@@
Shown in Fig. 4A is the soft robotic finger mounted on a three-axis motion platform for interactive deformation estimation. The test platform is operated manually to generate a set of contact configurations between the soft finger and the indenter. During the process, the in-finger camera streams real-time image data at a resolution of 640 × 480 pixels. Using an off-the-shelf ArUco detection library, the detected AMH rigid motion is fed into our implemented program for deformation estimation.
@@
A motion capture system (Mars2H by Nokov, Inc.) was used to track finger deformations through nine markers with an 8 mm radius. Among them, six markers were divided into three pairs, which were rigidly attached to the fingertip ($m_{5},m_{6}$), the first layer ($m_{3},m_{4}$), and the second layer ($m_{1},m_{2}$) of the soft finger, respectively. The other three markers were attached to the platform and used as the reference reading to align the motion capture system’s reference frame with the platform’s coordinate frame.
Figure 4: Estimated marker deformation obtained by proposed proprioceptive state estimation method. (A) Experimental setup, including the soft finger, embedded with an RGB camera, a manual three-axis motion test platform, and six motion capture markers $m_{1},m_{2},...,m_{6}$, rigidly attached to the soft finger. (B) The estimated position of the marker $x_{m_{k}}^{\prime}$ is calculated using the barycentric coordinate of the corresponding attached tetrahedron $t_{k}$, while the ground truth reading $x_{m_{k}}$ is obtained from the motion capture system. (C) The corresponding error for each marker’s three-dimensional deformation and total norm.
@@
The markers were attached to the soft finger with rigid links, as shown in Fig. 4B. We designed the connecting links to be three lengths to avoid occlusion during tracking. We assume each marker is rigidly attached to the nearest tetrahedron on the parameterized mesh model $\mathcal{M}$, representing the estimated marker location using barycentric coordinates of the corresponding tetrahedron element in the soft robotic finger’s deformed states:
@@
Due to the rigid connection assumption, the barycentric coordinates $\bm{\lambda}_{t_{k}}$ are constant during deformation. We solve the barycentric coordinates in Eq. (16) using the tetrahedron’s initial vertex position and corresponding tracked marker position without contact. The marker position prediction model is a linear combination of the deformed vertex position of the corresponding tetrahedron resulting from geometric optimization in Alg. 1 using calibrated barycentric coefficients. (See Movie S1 in the Supplementary Materials for a video demonstration.)
@@
We visualize the error distribution with 3k pairs of the six markers’ estimated and ground truth positions as illustrated in Fig. 4C. The norm of the six markers’ total error is within 3 mm, while error distribution along each axis is centered around the $(-2,2)$ mm range. As the marker prediction model in Eq. (16) comprises calibration and geometric optimization, the error distribution of six sparse markers may only partially validate the proposed method, leading to the next experiment.
4.1.3 Deformation Estimation using Touch Haptic Device
We designed another validation experiment using the pen-nib’s position of a haptic device (Touch by 3D Systems, Inc.) as ground truth measurement. As shown in Fig. 5A, an operator holding the pen-nib initiated contact at a random point on the soft robotic finger by pushing it five times. Fifty points were sampled, spreading over half of the soft robotic finger with recorded pen-nib position and the corresponding point of contact on the estimated deformation in the mesh model.